Can you show me Step-By-Step how to solve this problem?
Luanne bought 50 stamps, some costing 22 cents each, and the rest costing 14 cents each. She paid the postal clerk $10.04 for all the stamps. How many of each type of stamp did luanne buy?

Answer by scott8148(5891)

let x=(22 cent stamps), so 50-x=(14 cent stamps)

the number of stamps multiplied by their value gives the total cost, $10.04
... working in cents avoids problems with decimal points

22(x)+14(50-x)=1004 ... 22x+700-14x=1004 ... 8x=304 ... x=38

50-(38)=12